[Hallc_running] [New Logentry] RC daily update

hszumila at jlab.org hszumila at jlab.org
Wed Sep 26 08:35:02 EDT 2018

Logentry Text:
We are currently starting a sweep of Hall C for the following accesses this morning:
-HMS Q2 couldn't be turned on last evening due to a communication error with the magnet. We power cycled it, and it still didn't work (nor could be set manually at the power supply). Steve Lassiter is going in to look at the HMS Q2 power supply and SHMS encoder. 
-Hall C techs are going in for some other work.
-Dave Meekins is beginning target cool down and will also make an access. 

When beam comes back, this morning we will begin with getting the Hall C optics we want: smaller beam size at the target and high currents. Last night ops was able to send up to 50 uA CW to Hall C after dealing with issues from a few cavities found to be draining. They were also able to minimize the low energy tail in the 1st spreader. Once beam comes back, we will take a look at the beam profile in our Hall. We have a carbon target ion chamber calibration good up to 20 uA. 

As soon as we know the optics, we can begin to set up for the trigger and detector commissioning. I'll get into contact with the experts when I have a better idea of the time line. 

We will meet this afternoon at 4:30pm in the second floor counting house: https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and_daily_runplan_fall_2018


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3597964
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180926/95a19eee/attachment.html>

More information about the Hallc_running mailing list